# Kondo Screening and Indirect Magnetic Exchange through a Conventional Superconductor Studied by the Density-Matrix Renormalization Group

### v1

@@ VI.1 Exact-diagonalization results
For the $L=3$ model, Fig. 9 displays the boundaries between ground states with different total-spin quantum numbers. Figs. 10 and 11 show the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ and the local Kondo correlation $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle$, respectively. For strong $J$ and for $\Delta$ with $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta_{\rm c}(J)$, the ground state is a degenerate spin doublet ($S=\frac{1}{2}$), see blue region in Fig. 9. For large $\Delta\to\infty$, it is of the PKS form, Eq. (7), with correlations ${\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\to 0$ and ${\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\to-\frac{3}{8}$ (see Figs. 10 and 11). This deep PKS regime in the parameter space perfectly matches with the corresponding one for large $L$ ($L=83$, see Fig. 2). This is due to the fact that correlations are almost completely local for strong $J$ and for $\Delta\gg\delta$.
The orange-colored phase in Fig. 9 is an LKS phase with ${\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\to-\frac{3}{4}$ for $J\to\infty$, and the red-colored one an RKKY-singlet phase with ${\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\approx-\frac{3}{4}$ (see Figs. 10 and 11). In both cases, the ground state is a spin singlet. However, the eigenvalue of the $x$ component of the pseudo-spin, the pseudo-charge $q$, is different: We have $q=+\frac{1}{2}$ in the RKKY singlet and $q=-\frac{1}{2}$ in LKS ground state, while $q=0$ for the PKS state.
@@
Contrary, for $\Delta=0$ and for the small system with $L=3$, we have the situation of a “Kondo-vs.-RKKY quantum box”, as discussed in Refs. Thimm et al. (1999); Schwabe et al. (2012, 2015). Again, perturbation theory in $J$ is regular, but qualitatively different, since there is a finite “Kondo” term in the effective Hamiltonian
@@
see Ref. Schwabe et al. (2012) (supplemental material). The Kondo coupling $J_{\rm K}$ is positive (antiferromagnetic) and linear in $J$, while $J_{\rm RKKY}\propto J^{2}$, such that the finite-size Kondo effect “wins” over the RKKY coupling at sufficiently weak $J$. The mechanism for the Kondo coupling $J_{\rm K}$ in the toy model is routed in the $J=0$ electronic structure: There are three spin-degenerate conduction-electron eigenstates with one-particle energies $-\sqrt{2}t,0,\sqrt{2}t$. At half-filling, the one with lowest (highest) energy is fully occupied (unoccupied), while the delocalized and spin-degenerate “Fermi” orbital $\left|\rm F\right\rangle$ with zero one-particle energy is filled with either a $\sigma=\,\uparrow$ or a $\sigma=\,\downarrow$ electron. Hence, the $J=0$ many-body electronic ground state is degenerate (opposed to the case $\Delta>0$) and provides a single Kondo screening channel. For $d=2$ both impurity spins in fact couple to the electron spin ${\boldsymbol{s}}_{\rm F}=\frac{1}{2}\sum_{\sigma\sigma^{\prime}}c^{\dagger}_{F\sigma}{\boldsymbol{\tau}}_{\sigma\sigma^{\prime}}c_{F\sigma^{\prime}}$ of the Fermi orbital $\left|\rm F\right\rangle$. However, only one of the impurity spins can be screened, the other one must remain decoupled. This explains that the ground state is a total-spin doublet and has the form given by Eq. (7), i.e., a PKS state but with the difference that the Kondo screening of the impurity spin ${\boldsymbol{S}}_{m}$ is not local but achieved by the (same) delocalized Fermi orbital for both, $m=1$ and $m=2$. This implies that the according PKS doublet ground state takes the form
@@
Roughly, the finite-size Kondo effect dominates up to a critical local exchange $J_{\rm c}$, where the bulk Kondo temperature $T_{\rm K}\sim e^{-1/J}$ equals the finite-size gap $\delta$. At intermediate $J\gtrsim J_{\rm c}$, the RKKY interaction becomes dominant. In fact, this turns out as antiferromagnetic, $J_{\mathrm{RKKY}}=\frac{3J^{2}}{32\cdot\sqrt{2}}$. The RKKY interaction competes with the finite-size Kondo effect for weak $J$ and, for strong $J$, with local-Kondo-singlet formation, which leads to $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$. It is well known Schwabe et al. (2012), however, that the RKKY coupling becomes fully effective only for larger systems. This implies that, with increasing $J$, the impurity correlation decreases but remains positive (ferromagnetic) until it vanishes for $J\to\infty$ (see Fig.10).
@@ VI.2 Thermodynamical limit: extended discussion
On the strong-$J$ side of the phase diagram and for small $\Delta$, the impurity-spin correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$ and the local Kondo correlation $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to-\frac{3}{4}$ become identical in both, the doublet PKS and the singlet LKS phase (see Figs. 4 and 5 for $L=83$). The transition line $\Delta_{\rm c}^{\prime\prime}(J)$ in Fig. 2, which separates the $S=0$ LKS from the $S=\frac{1}{2}$ PKS phase at small $\Delta$, decreases with increasing $J$ and tends to $\Delta=0$ for $J\to\infty$. One possibility to approach this point in the phase diagram is via paths such that $\Delta>\Delta_{\rm c}^{\prime\prime}(J)$, i.e., staying in the LKS phase. In this case two fully localized Kondo singlets at sites $i_{1}$ and $i_{2}$ are formed and, furthermore, two BCS singlets with supports on the two (for $L\to\infty$) half-infinite chains $i=1,...,i_{1}-1$ and on $i=i_{2}+1,...,L$, and, finally, a completely localized pseudo-spin-$\eta_{ix}=\frac{1}{2}$ state $(|{\rm vac.}\rangle-|\uparrow\downarrow\rangle)/\sqrt{2}$. This means that we have a situation, where both impurity spins are screened separately. Another possibility is given by paths with $\Delta<\Delta_{\rm c}^{\prime\prime}(J)$, i.e., staying in the PKS phase. Since for $J\to\infty$ the impurity and the Kondo correlations functions become the same in PKS and the LKS phases, the physical situation does not change either, except for the point that the localized state at $i_{1}+1$ is a spin doublet $c^{\dagger}_{i_{1}+1\sigma}|{\rm vac.}\rangle$ ($\sigma=\uparrow,\downarrow$). The state is different from but smoothly connected to the state in the deep PKS limit at large $\Delta$ (with $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta_{\rm c}(J)$), where the doublet is formed by a single unscreened impurity spin. With decreasing $J$ and at $\Delta=0$, we expect that the impurity spins are still separately screened but that the two Kondo clouds extend spatially.
@@ VI.3 Self-consistency
Our numerical results are shown in Fig. 12 for comparison with the non-self-consistent results shown in Figs. 9, 10 and 11. With Fig. 12 we focus on the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$, which is most instructive for the qualitative discussion. There are four distinct phases with very different impurity-spin correlations.
@@
There is a small parameter region at intermediate $J$ and strong $|U|$ with weakly ferromagnetic correlations (see pink color in Fig. 12) and with quantum numbers $S=\frac{1}{2}$ and $q=0$. This is the spin-doublet PKS phase, which, at self-consistency, exhibits a spontaneous breaking of the spatial parity symmetry: We find $|\Delta_{1}|<|\Delta_{2}|<|\Delta_{3}|$ with a slightly negative $\Delta_{1}<0$ while $\Delta_{2},\Delta_{3}$ are almost equal and positive. In fact, at $i=1$, where the pairing strength $|\Delta_{1}|$ is small, the local Kondo correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{s}}_{i_{1}}\rangle$ is strongly antiferromagnetic while $\langle{\boldsymbol{S}}_{2}{\boldsymbol{s}}_{i_{2}}\rangle\approx 0$ since $\Delta_{i}$ has the maximum at $i=3$ (note $i_{1}=1$, $i_{2}=3$). So this is a PKS state, but compared to Eq. (7), the screening is selective and involves only a single impurity spin. Note that within the mean-field theory there is a second solution, which can be obtained by mirroring at $i=2$. In the phase diagram, this total-spin doublet PKS phase separates the RKKY-singlet from the LKS state. With increasing $|U|$, the PKS phase becomes narrower in the phase diagram Fig. 12 but its extension remains finite. So far, the phase diagram thus has the same topology as in the non-self-consistent case discussed in Sec. VI.1.
@@ VII Concluding remarks
A system of two quantum spins $\frac{1}{2}$ locally exchange coupled to a conventional s-wave superconductor features a competition between Kondo-singlet formation, BCS-singlet formation, and indirect magnetic exchange, which gives rise to a nontrivial phase diagram. Since the emergent nonlocal RKKY exchange between the impurity spins involves higher-energy excitations, this competition is expected to depend decisively on the system geometry. This dependence cannot be fully captured in approaches where the magnetic exchange is treated within mean-field approaches, such as dynamical mean-field theory Titvinidze et al. (2012), or where the RKKY coupling is introduced as an independent model parameter Žitko et al. (2010, 2011); Yao et al. (2014). We have studied the resulting phase diagram for a generic one-dimensional chain geometry, which is accessible to a DMRG approach that fully exploits the non-abelian spin-SU(2) symmetry.
@@
Our results for the one-dimensional model can be summarized as follows: First, the phase diagrams for impurities coupled to nearest or to next-nearest neighbor sites are fundamentally different. In the former case, there are smooth crossovers only with impurity-spin correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$, ranging from strongly antiferromagnetic for large pairing strengths $\Delta$ to almost vanishing correlations for strong $J$. Contrary, in the latter case, for an even distance $d=2$, we find discontinuous first-order (level-crossing) transitions between four phases: an RKKY-singlet and a local Kondo-singlet phase with impurity correlations similar as in the $d=1$ phase diagram, but in addition there is a partial Kondo screening (PKS) phase, a total-spin doublet, which separates the two singlet phases in the parameter space, and finally, at weak $J$ and small $\Delta$, an RKKY spin-triplet phase. As compared to previous NRG calculations Yao et al. (2014), the topology of the phase diagram is different. In particular, the PKS phase is stable in a rather wide range of exchange interactions $J$ at small $\Delta$, and appears to merge with the gapless $\Delta=0$.
@@
A finite gap, $\Delta>0$, is beneficial for the DMRG calculations because it leads to exponentially decaying RKKY and Kondo correlations. In fact, reaching the limit $\Delta\to 0$ is computationally increasingly costly for chains that are sufficiently long to keep finite-size artifacts under control. We have thus additionally considered a small system with $L=3$ sites only and two impurity spins at distance $d=2$, which can be treated by exact diagonalization and which gives complementary insights. Apart from the absence of an RKKY-triplet phase, which is well understood within RKKY perturbation theory, the ground-state phase diagram of this quantum-box model is qualitatively the same as the DMRG phase diagram for $d=2$, large $L$, and for $\Delta\gtrsim 0.1$. In the quantum-box model the $\Delta\to 0$ limit is easily accessible and exhibits a PKS doublet phase, which extends over the whole $J$ range. This supports the view that the same holds true for large $L$ in the $\Delta\to 0$ limit and, hence, that the phase boundary between the PKS and the RKKY (singlet or triplet) phase does not end in a critical point at a finite $J$ for $\Delta=0$, i.e., for the gapless system. We note that, at $\Delta=0$ and for small $L$, a re-entrant finite-size Kondo effect at weak $J$ on a linear-in-$J$ energy scale is present Thimm et al. (1999); Schwabe et al. (2012). It is known, however, that this is limited to smaller and smaller local exchange couplings $J$ with increasing $L$ and thus eventually becomes irrelevant in the thermodynamical limit $L\to\infty$.
@@
Using a Lanczos-transformation technique Büsser et al. (2013); Allerdt et al. (2015); Allerdt and Feiguin (2019), it is possible to map impurity-spin systems in rather arbitrary and higher-dimensional lattice geometries onto a one-dimensional chain model. This opens the perspective for further investigations of similar systems in various geometries, directly relevant to real materials and experimental studies, e.g., of magnetic atoms adsorbed on superconducting surfaces.
Another interesting avenue of further research is to study the effect of the spatial structure of a self-consistently determined pairing strength $\Delta_{i}$. Our results for the quantum-box model at $L=3$ indicate that a site-dependent $\Delta_{i}$ does not qualitatively change the phase-diagram topology but affects the parameter space, where the PKS phase is stable. Furthermore, a theoretical approach beyond a BCS-type mean-field theory for the conventional superconductor is interesting, see for example Ref. Pavešić et al. (2021).

### v2

@@ VI.1 Exact-diagonalization results
For the $L=3$ model, Fig. 8 displays the boundaries between ground states with different total-spin quantum numbers. Figs. 9 and 10 show the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ and the local Kondo correlation $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle$, respectively. For strong $J$ and for $\Delta$ with $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta_{\rm c}(J)$, the ground state is a degenerate spin doublet ($S=\frac{1}{2}$), see blue region in Fig. 8. For large $\Delta\to\infty$, it is of the PKS form, Eq. (7), with correlations ${\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\to 0$ and ${\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\to-\frac{3}{8}$ (see Figs. 9 and 10). This deep PKS regime in the parameter space perfectly matches with the corresponding one for large $L$ ($L=83$, see Fig. 2). This is due to the fact that correlations are almost completely local for strong $J$ and for $\Delta\gg\delta$.
The orange-colored phase in Fig. 8 is an LKS phase with ${\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\to-\frac{3}{4}$ for $J\to\infty$, and the red-colored one an RKKY-singlet phase with ${\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\approx-\frac{3}{4}$ (see Figs. 9 and 10). In both cases, the ground state is a spin singlet. However, the eigenvalue of the $x$ component of the pseudo-spin, the pseudo-charge $q$, is different: We have $q=+\frac{1}{2}$ in the RKKY singlet and $q=-\frac{1}{2}$ in LKS ground state, while $q=0$ for the PKS state.
@@
Contrary, for $\Delta=0$ and for the small system with $L=3$, we have the situation of a “Kondo-vs.-RKKY quantum box”, as discussed in Refs. [25, 36, 37]. Again, perturbation theory in $J$ is regular, but qualitatively different, since there is a finite “Kondo” term in the effective Hamiltonian
@@
see Ref. [36] (supplemental material). The Kondo coupling $J_{\rm K}$ is positive (antiferromagnetic) and linear in $J$, while $J_{\rm RKKY}\propto J^{2}$, such that the finite-size Kondo effect “wins” over the RKKY coupling at sufficiently weak $J$. The mechanism for the Kondo coupling $J_{\rm K}$ in the toy model is routed in the $J=0$ electronic structure: There are three spin-degenerate conduction-electron eigenstates with one-particle energies $-\sqrt{2}t,0,\sqrt{2}t$. At half-filling, the one with lowest (highest) energy is fully occupied (unoccupied), while the delocalized and spin-degenerate “Fermi” orbital $\left|\rm F\right\rangle$ with zero one-particle energy is filled with either a $\sigma=\,\uparrow$ or a $\sigma=\,\downarrow$ electron. Hence, the $J=0$ many-body electronic ground state is degenerate (opposed to the case $\Delta>0$) and provides a single Kondo screening channel. For $d=2$ both impurity spins in fact couple to the electron spin ${\boldsymbol{s}}_{\rm F}=\frac{1}{2}\sum_{\sigma\sigma^{\prime}}c^{\dagger}_{F\sigma}{\boldsymbol{\tau}}_{\sigma\sigma^{\prime}}c_{F\sigma^{\prime}}$ of the Fermi orbital $\left|\rm F\right\rangle$. However, only one of the impurity spins can be screened, the other one must remain decoupled. This explains that the ground state is a total-spin doublet and has the form given by Eq. (7), i.e., a PKS state but with the difference that the Kondo screening of the impurity spin ${\boldsymbol{S}}_{m}$ is not local but achieved by the (same) delocalized Fermi orbital for both, $m=1$ and $m=2$. This implies that the according PKS doublet ground state takes the form
@@
Roughly, the finite-size Kondo effect dominates up to a critical local exchange $J_{\rm c}$, where the bulk Kondo temperature $T_{\rm K}\sim e^{-1/J}$ equals the finite-size gap $\delta$. At intermediate $J\gtrsim J_{\rm c}$, the RKKY interaction becomes dominant. In fact, this turns out as antiferromagnetic, $J_{\mathrm{RKKY}}=\frac{3J^{2}}{32\cdot\sqrt{2}}$. The RKKY interaction competes with the finite-size Kondo effect for weak $J$ and, for strong $J$, with local-Kondo-singlet formation, which leads to $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$. It is well known [36], however, that the RKKY coupling becomes fully effective only for larger systems. This implies that, with increasing $J$, the impurity correlation decreases but remains positive (ferromagnetic) until it vanishes for $J\to\infty$ (see Fig.9).
@@ VI.2 Thermodynamical limit: extended discussion
On the strong-$J$ side of the phase diagram and for small $\Delta$, the impurity-spin correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle\to 0$ and the local Kondo correlation $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to-\frac{3}{4}$ become identical in both, the doublet PKS and the singlet LKS phase (see Fig. 4 for $L=83$). The transition line $\Delta_{\rm c}^{\prime\prime}(J)$ in Fig. 2, which separates the $S=0$ LKS from the $S=\frac{1}{2}$ PKS phase at small $\Delta$, decreases with increasing $J$ and tends to $\Delta=0$ for $J\to\infty$. One possibility to approach this point in the phase diagram is via paths such that $\Delta>\Delta_{\rm c}^{\prime\prime}(J)$, i.e., staying in the LKS phase. In this case two fully localized Kondo singlets at sites $i_{1}$ and $i_{2}$ are formed and, furthermore, two BCS singlets with supports on the two (for $L\to\infty$) half-infinite chains $i=1,...,i_{1}-1$ and on $i=i_{2}+1,...,L$, and, finally, a completely localized pseudo-spin-$\eta_{ix}=\frac{1}{2}$ state $(|{\rm vac.}\rangle-|\uparrow\downarrow\rangle)/\sqrt{2}$. This means that we have a situation, where both impurity spins are screened separately. Another possibility is given by paths with $\Delta<\Delta_{\rm c}^{\prime\prime}(J)$, i.e., staying in the PKS phase. Since for $J\to\infty$ the impurity and the Kondo correlation functions become the same in PKS and the LKS phases, the physical situation does not change either, except for the point that the localized state at $i_{1}+1$ is a spin doublet $c^{\dagger}_{i_{1}+1\sigma}|{\rm vac.}\rangle$ ($\sigma=\uparrow,\downarrow$). The state is different from but smoothly connected to the state in the deep PKS limit at large $\Delta$ (with $\Delta^{\prime}_{\rm c}(J)<\Delta<\Delta_{\rm c}(J)$), where the doublet is formed by a single unscreened impurity spin. With decreasing $J$ and at $\Delta=0$, we expect that the impurity spins are still separately screened but that the two Kondo clouds extend spatially.
@@ VI.3 Self-consistency
Our numerical results are shown in Fig. 11 for comparison with the non-self-consistent results shown in Figs. 8, 9 and 10. With Fig. 11 we focus on the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$, which is most instructive for the qualitative discussion. There are four distinct phases with very different impurity-spin correlations.
@@
There is a small parameter region at intermediate $J$ and strong $|U|$ with weakly ferromagnetic correlations (see pink color in Fig. 11) and with quantum numbers $S=\frac{1}{2}$ and $q=0$. This is the spin-doublet PKS phase, which, at self-consistency, exhibits a spontaneous breaking of the spatial parity symmetry: We find $|\Delta_{1}|<|\Delta_{2}|<|\Delta_{3}|$ with a slightly negative $\Delta_{1}<0$ while $\Delta_{2},\Delta_{3}$ are almost equal and positive. In fact, at $i=1$, where the pairing strength $|\Delta_{1}|$ is small, the local Kondo correlation function $\langle{\boldsymbol{S}}_{1}{\boldsymbol{s}}_{i_{1}}\rangle$ is strongly antiferromagnetic while $\langle{\boldsymbol{S}}_{2}{\boldsymbol{s}}_{i_{2}}\rangle\approx 0$ since $\Delta_{i}$ has the maximum at $i=3$ (note $i_{1}=1$, $i_{2}=3$). So this is a PKS state, but compared to Eq. (7), the screening is selective and involves only a single impurity spin. Note that within the mean-field theory there is a second solution, which can be obtained by mirroring at $i=2$. In the phase diagram, this total-spin doublet PKS phase separates the RKKY-singlet from the LKS state. With increasing $|U|$, the PKS phase becomes narrower in the phase diagram Fig. 11 but its extension remains finite. So far, the phase diagram thus has the same topology as in the non-self-consistent case discussed in Sec. VI.1.
@@ VII Concluding remarks
A system of two quantum spins $\frac{1}{2}$ locally exchange coupled to a conventional s-wave superconductor features a competition between Kondo-singlet formation, BCS-singlet formation, and indirect magnetic exchange, which gives rise to a nontrivial phase diagram. Since the emergent nonlocal RKKY exchange between the impurity spins involves higher-energy excitations, this competition is expected to depend decisively on the system geometry. This dependence cannot be fully captured in approaches where the magnetic exchange is treated within mean-field approaches, such as dynamical mean-field theory [16], or where the RKKY coupling is introduced as an independent model parameter [68, 69, 70]. We have studied the resulting phase diagram for a generic one-dimensional chain geometry, which is accessible to a DMRG approach that fully exploits the non-abelian spin-SU(2) symmetry.
@@
Our results for the one-dimensional model can be summarized as follows: First, the phase diagrams for impurities coupled to nearest or to next-nearest neighbor sites are fundamentally different. In the former case, there are smooth crossovers only with impurity-spin correlations $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$, ranging from strongly antiferromagnetic for large pairing strengths $\Delta$ to almost vanishing correlations for strong $J$. Contrary, in the latter case, for an even distance $d=2$, we find discontinuous first-order (level-crossing) transitions between four phases: an RKKY-singlet and a local Kondo-singlet phase with impurity correlations similar as in the $d=1$ phase diagram, but in addition there is a partial Kondo screening (PKS) phase, a total-spin doublet, which separates the two singlet phases in the parameter space, and finally, at weak $J$ and small $\Delta$, an RKKY spin-triplet phase. As compared to previous NRG calculations [70], the topology of the phase diagram is different. In particular, the PKS phase is stable in a rather wide range of exchange interactions $J$ at small $\Delta$, and appears to merge with the gapless $\Delta=0$.
@@
A finite gap, $\Delta>0$, is beneficial for the DMRG calculations because it leads to exponentially decaying RKKY and Kondo correlations at large distances. In fact, reaching the limit $\Delta\to 0$ is computationally increasingly costly for chains that are sufficiently long to keep finite-size artifacts under control. We have thus additionally considered a small system with $L=3$ sites only and two impurity spins at distance $d=2$, which can be treated by exact diagonalization and which gives complementary insights. Apart from the absence of an RKKY-triplet phase, which is well understood within RKKY perturbation theory, the ground-state phase diagram of this quantum-box model is qualitatively the same as the DMRG phase diagram for $d=2$, large $L$, and for $\Delta\gtrsim 0.1$. In the quantum-box model the $\Delta\to 0$ limit is easily accessible and exhibits a PKS doublet phase, which extends over the whole $J$ range. This supports the view that the same holds true for large $L$ in the $\Delta\to 0$ limit and, hence, that the phase boundary between the PKS and the RKKY (singlet or triplet) phase does not end in a critical point at a finite $J$ for $\Delta=0$, i.e., for the gapless system. We note that, at $\Delta=0$ and for small $L$, a re-entrant finite-size Kondo effect at weak $J$ on a linear-in-$J$ energy scale is present [25, 36]. It is known, however, that this is limited to smaller and smaller local exchange couplings $J$ with increasing $L$ and thus eventually becomes irrelevant in the thermodynamical limit $L\to\infty$.
@@
Using a Lanczos-transformation technique [92, 19, 21], it is possible to map impurity-spin systems in rather arbitrary and higher-dimensional lattice geometries onto a one-dimensional chain model. This opens the perspective for further investigations of similar systems in various geometries, directly relevant to real materials and experimental studies, e.g., of magnetic atoms adsorbed on superconducting surfaces. In this context, it will be important to additionally consider anisotropic interactions and to extend the studies to multi-orbital models.
Another interesting avenue of further research is to study the effect of the spatial structure of a self-consistently determined pairing strength $\Delta_{i}$. Our results for the quantum-box model at $L=3$ indicate that a site-dependent $\Delta_{i}$ does not qualitatively change the phase-diagram topology but affects the parameter space, where the PKS phase is stable. In addition, the computation and discussion of the nonlocal pairing correlation function $\langle c_{i\downarrow}c_{i\uparrow}\,c_{j\uparrow}^{\dagger}c_{j\downarrow}^{\dagger}\rangle$ is worthwhile, especially for the inhomogeneous case with site-dependent $\Delta_{i}$. Furthermore, a theoretical approach beyond a BCS-type mean-field theory for the conventional superconductor is interesting, see for example Ref. [93].